The following information is provided for Southall Company: Sales revenue Variable manufacturing costs Fixed manufacturing coMultiple Choice o $30,000 $45,000 0 $67,500 0 $17,500 $17,500
Answer #1

Total variable cost=Variable manufacturing costs+Variable selling and administrative costs

=(42500+15000)=$57500

Contribution margin=Sales-Variable cost

=125,000-57500

=$67500